OpenStudy (anonymous):

F(x) varies directly with x². If f(x) = 12 when x = 2, write the equation for the variation and graph it showing at least three points. Draw a chart to show at least three ordered pairs that are on the graph.

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

\[f(x)=kx^2\\f(2)=k\times 2^2=12\\ 4k=12\\ k=3\] so \[f(x)=3x^2\)

OpenStudy (misty1212):

oops \[f(x)=3x^2\] plot that one